OpenStudy (anonymous):

Percents are ratios that compare numbers to _________. A. 1/2 B. 0 C. 100 D. 1

OpenStudy (anonymous):

so these fractions are supposed to be compared right?

OpenStudy (anonymous):

im not sure exactly

OpenStudy (anonymous):

i think its a or c from somewhere i seen

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }=0.5= 50 \] %

OpenStudy (anonymous):

A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Write 80% as a decimal and fraction (in lowest terms). A. 80 and 80/100 B. 0.8 and 8/10 C. 8 and 80/100 D. 0.8 and 4/5

OpenStudy (anonymous):

\[80(percent)=\frac{ 80 }{ 100 }= 0.8\] okay?

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

yeah it just gets reduced

OpenStudy (anonymous):

no it's D

OpenStudy (anonymous):

because it got reduced.

OpenStudy (anonymous):

yes \[\frac{ 80 }{ 100 }=\frac{ 4 }{ 5 }\]